# FLARE: Fast Low-rank Attention Routing Engine

### v1

@@ Abstract
The quadratic complexity of self-attention limits its applicability and scalability on large unstructured meshes. We introduce Fast Low-rank Attention Routing Engine (FLARE), a linear complexity self-attention mechanism that routes attention through fixed-length latent sequences. Each attention head performs global communication among $N$ tokens by projecting the input sequence onto a fixed length latent sequence of $M\ll N$ tokens using learnable query tokens. By routing attention through a bottleneck sequence, FLARE learns a low-rank form of attention that can be applied at $\mathcal{O}(NM)$ cost. FLARE not only scales to unprecedented problem sizes, but also delivers superior accuracy compared to state-of-the-art neural PDE surrogates across diverse benchmarks. We also release a new additive manufacturing dataset to spur further research. Our code is available at https://github.com/vpuri3/FLARE.py.
@@ 1 Introduction
In the context of learning partial differential equation (PDE) surrogates, i.e., regressing continuous fields over point clouds/meshes, each point in the 3D point cloud is treated as a token, with associated features that encode geometric and physical quantities such as coordinates, normals, and material properties. Unlike sequential tokens in language models, these tokens are spatially distributed with no inherent ordering to them.
Several works have proposed approximations to the full self-attention mechanism to mitigate the cost. Perceiver and PerceiverIO (Jaegle et al., 2021b; a) introduced cross-attention between the input sequence and a length $M$ learnable latent sequence. In this mechanism, the $M\times N$ attention weights are used to project the input sequence to the latent sequence, followed by self-attention blocks in the latent space. Perceiver thus decouples the length of the input sequence from the computational complexity by avoiding dense self-attention on the inputs. However, the latent bottleneck in PerceiverIO can limit accuracy as the model may discard fine-grained features if the number of latent tokens is too low.
@@
  • Note: *Physics Attention in Transolver uses same projection weights for all heads whereas PerceiverIO and FLARE use cross-attention projection where each head learns a distinct slice of the latent query tokens.
We propose Fast Low-rank Attention Routing Engine (FLARE), a simple yet powerful mechanism designed to break the scalability barrier in PDE surrogate learning. FLARE is built on the argument that projecting input sequences onto shorter latent sequences, and then unprojecting to the original sequence length, is equivalent to constructing a low-rank form of attention with rank at most equal to the number of latent tokens (see Figure 1).
@@
  • Linear complexity token mixing. FLARE is an efficient self-attention mechanism designed to learn on long sequences such as point clouds. By replacing full self-attention with low-rank projections and reconstructions, FLARE achieves linear complexity in the number of points.
@@
  • Benchmark dataset of additive manufacturing simulations. We release a large-scale, high-resolution dataset for residual displacement prediction in metal additive manufacturing to encourage further research in scalable PDE surrogate modeling.
@@ 2 Related work
Several methods have been proposed to address the quadratic complexity of the attention mechanism. In (Wang et al., 2020), attention matrices are shown to be low-rank. Their proposed model, called Linformer, employs learned linear mappings to project the input key and value sequences to a shorter sequence length, resulting in a linear complexity form of attention. A drawback of this approach is that it cannot directly handle variable input sequence lengths. Another efficient attention method, called multi-head latent attention (Liu et al., 2024), jointly compresses key and value tokens to reduce the memory footprint of language models during inference. However, this does not help with the quadratic bottleneck in self-attention. Low-Rank Adaptation (LoRA) (Hu et al., 2022) also leverages low-rank projections, though primarily for efficient fine-tuning rather than architectural design or modeling capacity, as we explore in this work.
While these efficient attention methods are general frameworks, none specifically targets PDE surrogate modeling. Our proposed FLARE method leverages low-rankness, is sequence-length independent, and achieves competitive performance on PDE benchmarks while delivering a substantial speedup compared to other existing methods.
@@ 3.1 Preliminary: Multi-Head Self-Attention
The greatest cost in multi-head self-attention is the call to SDPA which is $\mathcal{O}(N^{2})$ in complexity as both the query and key sequences have $N$ tokens. Fortunately, SDPA is a fundamental algorithm in modern deep learning, and its GPU optimized multi-head implementations are available in PyTorch (Paszke, 2019).
@@ 3.2 FLARE: Fast Low-rank Attention Routing Engine
Formally, we define a learnable query matrix $Q\in\mathbb{R}^{M\times C}$, where each row corresponds to a latent token. The key and value matrices, $K,V\in\mathbb{R}^{N\times C}$, are obtained by applying deep residual multi-layer perceptrons (MLPs) detailed in Appendix A to the input $X$. Compared to just a linear layer, these allow the model to learn higher-order feature interactions and deeper nonlinear transformations. Refer to Section 4.3 for ablation studies.
@@
where $Y_{h}\in\mathbb{R}^{N\times D}$ is the output for each head. Similar to multi-head self attention, the outputs from all heads are concatenated along the feature dimension and passed through a final linear projection to mix information across heads. As the query matrix has only $M$ tokens, the cost of SDPA calls in Eq. 5 and Eq. 6 is $\mathcal{O}(NM)$. PyTorch code for the multi-head implementation is presented in Figure 3. Note that we use a scaling factor $s=1$ instead of $\sqrt{D}$ in typical transformers (Vaswani et al., 2017)) in SDPA. This modification is explained in Section 4.3.
@@
Note that $\mathrm{softmax}$ is taken along the row dimension. We define
@@
as the dense global communication matrix with rank at most $M$. This low-rank structure, illustrated in Figure 1, enables efficient all-to-all communication without explicitly forming $W_{h}$; instead, $W_{\text{encode},h}$ and $W_{\text{decode},h}$ are applied sequentially, resulting in an overall cost of $\mathcal{O}(MN)$ per head. Finally, note that $W_{\text{encode},h}$ and $W_{\text{decode},h}$ are adjoints of one another up to diagonal scaling. This results in a near symmetric attention pattern.
@@ 3.3 Spectral analysis
The matrix $W_{h}$ in Eq. 9 represents the attention weights between $N$ tokens, where $W_{h}]_{ij}$ quantifies how much token $j$ communicates to token $i$ within head $h$. Since $W_{h}$ has rank at most $M$, the model can capture at most $M$ independent global communication patterns per head. The eigenvalues of $W$ indicate the relative importance or energy of each latent dimension in forming the attention matrix $W$. We devise an algorithm to obtain the eigen decomposition of $W$ in $\mathcal{O}(M^{3}+M^{2}N)$ time compared to $\mathcal{O}(N^{3})$ for a dense communication matrix. The algorithm is predicated on computing the eigenspectra of an $M\times M$ matrix $JJ^{T}$ where $J\in\mathbb{R}^{M\times N}$ is chosen such that $J^{T}J$ is similar to $W$. The algorithm is detailed in Section B.1, and summarized in Algorithm 1. Figure 8 presents the $M$ nonzero eigenvalues of $W_{h}$ for an FLARE model trained on the elasticity benchmark problem with 972 points per input. The distinct spectra of the heads indicates that each head learns distinct attention patterns.
@@ 4.1 PDE surrogate benchmarks
We consider a diverse set of benchmark datasets (summarized in Appendix Table 3) for regressing PDE solutions on point clouds. The test suite includes 2D and 3D benchmark datasets on structured and unstructured meshes with up to 50,000 points used for training and evaluation. Note that FLARE is agnostic to the underlying mesh structure as the model only considers the input point cloud. The 2D benchmark problems of Elasticity, Darcy, Airfoil, and Pipe (Li et al., 2020; 2023b) collectively simulate a wide range of physical phenomena. Among the 3D benchmarks, the DrivAerML (Ashton et al., 2024) benchmark problem contains automotive aerodynamic simulations. More details about the datasets are provided in Appendix C.2.
Finally, we propose a 3D field prediction benchmark derived from numerical simulations of the laser powder bed fusion (LPBF) additive manufacturing process. The LPBF benchmark task entails predicting the deformation field over a diverse set of 3D printed parts with up to 50,000 grid points. More details on this proposed benchmark problem are presented in Appendix E.
@@
We compare FLARE with state-of-the-art PDE surrogate models. The comparison suite includes generic attention-based models such as vanilla transformer (Vaswani et al., 2017), PerceiverIO (Jaegle et al., 2021a), attention-based PDE surrogate models such as Transolver (Wu et al., 2024), LNO (Wang & Wang, 2024a)), and a neural operator model GNOT (Hao et al., 2023). We do not include any graph-based model in our comparison suite, as the graph connectivity information is not available for most problems. Transolver++ (Luo et al., 2025) would have been a valuable addition to the comparison suite, but its code was not publicly available at the time of writing.
We follow the experimental setup of Transolver as it is the preeminent surrogate model and attempt to match its parameter count. Note that Transolver can be instantiated in two configurations: without convolution, where point-to-point communication relies solely on physics attention, and with convolution, where convolution layers are added to inject information from neighboring points when the input grid is structured. We evaluate these two configurations separately to isolate the impact of convolution versus physics attention. In our model, we choose not to employ any convolution layers and rely entirely on FLARE for token mixing.
Finally, we note that the PDE problems in this section are relatively small (up 50,000 points), and opt to train all models in full-precision (FP32). As shown in Figure 9, the vanilla transformer becomes drastically slower in comparison to FLARE and Transolver for on large point clouds. As such, we only evaluate the vanilla transformer on problems with up to $\sim$10,000 points.
@@ Discussion.
Although the vanilla transformer is extremely effective on small-scale PDE problems, it becomes extremely slow on large point clouds due to its quadratic cost as illustrated in Figure 2. On the contrary, PerceiverIO (with only a single encoding and decoding step) performs poorly even with $M=1,024$ latent tokens and $B=8$ latent self-attention blocks. This validates our hypothesis that multiple latent self-attention operations can be unnecessary and potentially suboptimal so long as the projections are sufficiently expressive. This is because information loss during projection is not recoverable via latent self-attention alone. Instead, performing multiple (head-wise) parallel projections and reconstructions directly between the input and latent sequences preserves expressivity while simplifying the architecture.
@@ 4.3 Model analysis and ablations
Although vanilla self-attention has the lowest memory cost thanks to the flash-attention algorithm, which eliminates the need to materialize the score matrices ($Q_{h}\cdot K_{h}^{T}$), its compute time still scales poorly with the sequence length. In contrast, the compute time for FLARE exhibits excellent scaling with sequence length. Its memory requirement is marginally greater than vanilla attention due to the presence of deep residual networks for key/value projections, and due to the need to materialize $Z_{h}$, the latent sequence of $M$ tokens. As these costs are marginal compared to the SDPA operation, the curves for different $M$ values of FLARE are somewhat overlapping. Finally, the compute time for Physics Attention of Transolver (Wu et al., 2024) exhibits somewhat good scaling. However, its memory cost and compute time blow up for large slice counts due to the need for materializing the projection matrices.
ResMLP depth: key/value projections.
ResMLP depth: feedforward block.
Number of heads.
A key hypothesis behind our design is employing multiple simultaneous (head-wise parallel) low-rank projections to collectively approximate a full attention map. Thus, for a fixed total feature dimension ($C$), using a smaller head dimension ($D$) should improve performance by allowing more such parallel projections. To test this, we evaluate FLARE on the elasticity dataset with varying head counts while keeping the total feature dimension $C=64$ and number of blocks $B=8$ fixed. As shown in Figure 6 (left), FLARE achieves the lowest test error when $D=4$ or $D=8$, in contrast to standard transformers which often use D=16D=163232. Notably, because our head dimensions are smaller, we use a scaling factor of $1$ instead of the typical $\sqrt{D}$ employed in transformers (Vaswani et al., 2017), as the dot product magnitudes are sufficiently small without additional scaling.
@@
Figure 6 (right) presents the test relative error of FLARE on the elasticity benchmark dataset as a function of the number of blocks $(B)$ for different latent sequence length $(M)$. Figure 4 (left) presents the same for the DrivAerML dataset with one million points per geometry. In both cases, we note the favorable trend that relative error consistently decreases as we increase the number of blocks. Similarly, we observe that the relative error generally decreases with $M$, though the trend is not strictly monotonic. Critically, we note in Figure 4 (right) that increasing $M$ does not come at the cost of greater memory requirements.
Figure 6: (Left) Effect of head dimension ($D$) on test accuracy. FLARE is designed to work optimally for D=4D=488. (Right) Effect of number of blocks ($B$) on test accuracy. Both experiments are conducted on the elasticity benchmark problem with $C=64$ feature dimension.

### v2

@@ Abstract
The quadratic complexity of self-attention limits its applicability and scalability on large unstructured meshes. We introduce Fast Low-rank Attention Routing Engine (FLARE), a linear complexity self-attention mechanism that routes attention through fixed-length latent sequences. Each attention head performs global communication among $N$ tokens by projecting the input sequence onto a fixed length latent sequence of $M\ll N$ tokens using learnable query tokens. Routing information through this bottleneck constrains the interaction to have rank at most $M$: the resulting operation is equivalent to an attention matrix of rank at most $M$, realized at $\mathcal{O}(NM)$ cost. This formulation allows users to flexibly adjust the effective rank of attention –and thus the balance between accuracy and efficiency– by tuning $M$ to their resource and performance requirements. FLARE not only scales to unprecedented problem sizes, but also delivers superior accuracy compared to state-of-the-art neural PDE surrogates across diverse benchmarks. We also introduce a new additive manufacturing dataset to spur further research. Our code is available at https://github.com/vpuri3/FLARE.py.
@@ 1 Introduction
In the context of learning partial differential equation (PDE) surrogates, i.e., regressing continuous fields over point clouds/meshes, each point in the 3D point cloud is treated as a token, with associated features that encode geometric and physical quantities such as coordinates, normals, and material properties.
Several works have proposed approximations to the full self-attention mechanism to mitigate the cost. Perceiver and PerceiverIO (Jaegle et al., 2021b; a) introduced cross-attention between the input sequence and a length $M$ learnable latent sequence. In this mechanism, the $M\times N$ attention weights are used to project the input sequence to the latent sequence, followed by self-attention blocks in the latent space. This decouples the length of the input sequence from the computational complexity by avoiding dense self-attention on the inputs. However, the latent bottleneck in PerceiverIO can limit accuracy as the model may discard fine-grained features if the number of latent tokens is too low.
@@
  • *Physics Attention in Transolver uses same projection weights for all heads whereas PerceiverIO and FLARE use cross-attention projection where each head learns a distinct slice of the latent query tokens.
We propose Fast Low-rank Attention Routing Engine (FLARE), a simple yet effective mechanism designed to break the scalability barrier in PDE surrogate learning. FLARE is built on the argument that projecting input sequences onto shorter latent sequences, and then unprojecting to the original sequence length, is equivalent to constructing a low-rank form of attention with rank at most equal to the number of latent tokens (see Figure 1).
@@
  • Linear complexity token mixing. FLARE is self-attention mechanism designed to learn on long sequences such as point clouds. By replacing full self-attention with low-rank projections and reconstructions, FLARE achieves linear complexity in the number of points.
@@
  • Benchmark dataset of additive manufacturing simulations. We release a large-scale, high-resolution dataset for residual displacement prediction in metal additive manufacturing to encourage further research in scalable PDE surrogate modeling. The dataset is available at https://huggingface.co/datasets/vedantpuri/LPBF_FLARE.
@@ 2 Related work
Several methods have been proposed to address the quadratic complexity of the attention mechanism. In (Wang et al., 2020), attention matrices are shown to be low-rank. Their proposed model, called Linformer, employs learned linear mappings to project the input key and value sequences to a shorter sequence length, resulting in a linear complexity form of attention. A drawback of this approach is that it cannot directly handle variable input sequence lengths. Other low-rank approaches include Reformer (Kitaev et al., 2020), which uses locality-sensitive hashing for efficient approximate attention, Nyströmformer (Xiong et al., 2021), which approximates self-attention using the Nyström method, and Scatterbrain (Chen et al., 2021), which combines sparse and low-rank methods for further gains. Another efficient attention method, called multi-head latent attention (Liu et al., 2024), jointly compresses key and value tokens to reduce the memory footprint of language models during inference. However, this does not help with the quadratic bottleneck in self-attention. Low-Rank Adaptation (LoRA) (Hu et al., 2022) also leverages low-rank projections, though primarily for efficient fine-tuning rather than architectural design or modeling capacity, as we explore in this work.
@@ 3.1 Preliminary: Multi-Head Self-Attention
The greatest cost in multi-head self-attention is the call to SDPA which is $\mathcal{O}(N^{2})$ in time and memory complexity. This is because $Q_{h}\cdot K_{h}^{T}\in\mathbb{R}^{N\times N}$ requires $\mathcal{O}(N^{2})$ storage and softmax, matrix-vector product with $V_{h}$ takes $\mathcal{O}(N^{2})$ operations. Fortunately, GPU optimized multi-head implementations of SDPA are available in PyTorch (Paszke, 2019).
@@ 3.2 FLARE: Fast Low-rank Attention Routing Engine
Formally, we define a learnable query matrix $Q\in\mathbb{R}^{M\times C}$, where each row corresponds to a latent token. The key and value matrices, $K,V\in\mathbb{R}^{N\times C}$, are obtained by applying deep residual multi-layer perceptrons (MLPs) detailed in Appendix A to the input $X$. Compared to just a linear layer, these allow the model to learn higher-order feature interactions and deeper nonlinear transformations. Refer to Appendix F for ablation studies.
@@
where $Y_{h}\in\mathbb{R}^{N\times D}$ is the output for each head. Similar to multi-head self attention, the outputs from all heads are concatenated along the feature dimension and passed through a final linear projection to mix information across heads. As the query matrix has only $M$ tokens, the cost of SDPA calls in Eq. 5 and Eq. 6 is $\mathcal{O}(NM)$. PyTorch code for the multi-head implementation is presented in Figure 3. Note that we use a scaling factor $s=1$ instead of $\sqrt{D}$ in typical transformers (Vaswani et al., 2017)) in SDPA. This modification is explained in Appendix F.
@@
Note that $\mathrm{softmax}$ is taken along the row-dimension. We define
@@
as the dense global communication matrix with rank at most $M$. This low-rank structure, illustrated in Figure 1, enables efficient all-to-all communication without explicitly forming $W_{h}$; instead, $W_{\text{encode},h}$ and $W_{\text{decode},h}$ are applied sequentially, resulting in an overall cost of $\mathcal{O}(MN)$ per head.
Discussion on the design principles of FLARE.
We defer to Appendix E a deeper analysis of how FLARE’s latent tokens enable gather–scatter communication (acting as selective pooling hubs and broadcasters), why the symmetry between encoding and decoding operators promotes stable information flow, and how fixing latent queries highlights a tradeoff between query dynamics and the need for expressive key/value projections. Together, these perspectives clarify the structural principles that underlie FLARE’s efficiency and expressivity.
@@ 3.3 Spectral analysis
The matrix $W_{h}$ in Eq. 9 represents the attention weights between $N$ tokens, where $W_{h}]_{ij}$ quantifies how much token $j$ communicates to token $i$ within head $h$. Since $W_{h}$ has rank at most $M$, the model can capture at most $M$ independent global communication patterns per head. The eigenvalues of $W$ indicate the relative importance or energy of each latent dimension in forming the attention matrix $W$. We apply an algorithm to obtain the eigen decomposition of $W$ in $\mathcal{O}(M^{3}+M^{2}N)$ time compared to $\mathcal{O}(N^{3})$ for a dense communication matrix. The algorithm is predicated on computing the eigenspectra of an $M\times M$ matrix $JJ^{T}$ where $J\in\mathbb{R}^{M\times N}$ is chosen such that $J^{T}J$ is similar to $W$. The algorithm is detailed in Section B.1, and summarized in Algorithm 1. Figure 7 presents the $M$ nonzero eigenvalues of $W_{h}$ for an FLARE model trained on the elasticity benchmark problem with 972 points per input. The distinct spectra of the heads indicates that each head learns distinct attention patterns.
@@ 4.1 PDE surrogate benchmarks
We consider a diverse set of benchmark datasets (Table 3) for regressing PDE solutions on point clouds spanning structured and unstructured grids with up to 50,000 points. Note that FLARE is mesh-agnostic, and operates solely on the input point cloud. The 2D elasticity, darcy, airfoil, and pipe benchmarks (Li et al., 2020; 2023b) cover a wide range of physical phenomena, and the 3D DrivAerML benchmark (Ashton et al., 2024) provides automotive aerodynamic simulations. Additional details of the dataset are presented in Section C.2. We also introduce a 3D field-prediction benchmark derived from laser powder bed fusion (LPBF) simulations, with diverse 3D-printed parts containing up to 50,000 grid points (see Appendix G).
@@
We compare FLARE with state-of-the-art PDE surrogates: generic attention models (vanilla Transformer (Vaswani et al., 2017), PerceiverIO (Jaegle et al., 2021a)); attention-based PDE surrogates (Transolver (Wu et al., 2024), LNO (Wang & Wang, 2024a)); and the neural operator GNOT (Hao et al., 2023). We exclude graph-based models because graph connectivity is unavailable for most problems. Transolver++ (Luo et al., 2025) is a concurrent work; although we tested the recently released official implementation with reported hyperparameters, preliminary results did not match those reported in the paper, so we do not include it in our comparisons. We follow the experimental setup of Transolver as it is the preeminent surrogate model and attempt to match its parameter count. Note that Transolver can be instantiated in two configurations: without convolution, where point-to-point communication relies solely on physics attention, and with convolution, where convolution layers are added to inject information from neighboring points when the input grid is structured. We evaluate these two configurations separately to isolate the impact of convolution versus physics attention. In our model, we choose not to employ any convolution layers and rely entirely on FLARE for token mixing. Because these PDE problems are relatively small (up to 50,000 points), we train all models in FP32. As shown in Figure 8, the vanilla Transformer is drastically slower than FLARE and Transolver on large point clouds; accordingly, we evaluate it only on problems up to $\sim 10{,}000$ points.
@@ Discussion.
Although the vanilla transformer is extremely effective on small-scale PDE problems, it becomes prohibitively slow on large point clouds due to its quadratic cost as illustrated in Figure 2. On the contrary, PerceiverIO (with only a single encoding and decoding step) performs poorly even with $M=1,024$ latent tokens and $B=8$ latent self-attention blocks. This validates our hypothesis that multiple latent self-attention operations can be unnecessary and potentially suboptimal so long as the projections are sufficiently expressive. This is because information loss during projection is not recoverable via latent self-attention alone. Instead, performing multiple (head-wise) parallel projections and reconstructions directly between the input and latent sequences preserves expressivity while simplifying the architecture.
@@ 4.3 Model analysis and ablations
In this section, we scale FLARE to examine how expressivity trades off against time and memory complexity. Complementary ablations in Appendix F analyze architectural choices such as the depth of residual MLPs in key/value projections, the design of the feedforward block, and the effect of head dimension on accuracy.
@@
Although vanilla self-attention has the lowest memory cost thanks to the flash-attention algorithm, which eliminates the need to materialize the score matrices ($Q_{h}\cdot K_{h}^{T}$), its compute time still scales poorly with the sequence length. In contrast, the compute time for FLARE exhibits strong scaling with sequence length. Its memory requirement is marginally greater than vanilla attention due to the presence of deep residual networks for key/value projections, and due to the need to materialize $Z_{h}$, the latent sequence of $M$ tokens. As these costs are marginal compared to the SDPA operation, the curves for different $M$ values of FLARE are somewhat overlapping. Finally, the compute time for Physics Attention of Transolver (Wu et al., 2024) exhibits somewhat good scaling. However, its memory cost and compute time blow up for large slice counts due to the need for materializing the projection matrices.
Figure 5: Effect of number of blocks ($B$) and number of latent tokens ($M$) on test accuracy on the elasticity (left) and darcy (right) test cases.
@@
Figure 5 presents the test relative error of FLARE on the elasticity (left) and darcy (right) benchmark datasets as a function of the number of blocks $(B)$ and the number of latent tokens $(M)$. Figure 4 (left) presents the same for the DrivAerML dataset with one million points per geometry. In all cases, we note the favorable trend that relative error consistently decreases as we increase the number of blocks. Similarly, we observe that the relative error generally decreases with $M$, though the trend is not strictly monotonic. In the elasticity problem, improvements with rank diminish rapidly, indicating that global communication in that problem is fundamentally low-rank. On the other hand, increasing $M$ monotonically increases performance on the darcy problem, indicating that the problem is rank-limited. This also explains why vanilla transformer with a full-rank attention pattern outperforms rank-deficient FLARE on the darcy problem. However, the accuracy gain comes at the cost of greater latency as the vanilla transformer is $\sim 5\times$ slower than FLARE on the Darcy problem. Figure 4 indicates that time per epoch (middle) and memory (right) scaling of FLARE with $B$ and $M$. Here, increasing $M$ leads to increased latency, and that increasing $M$ does not come at the cost of greater memory requirements.